Simon Teguh is considering investing in a vending machine operation involving 20 vending machines located in various plants around the city. The machine manufacturer reports that similar vending machine routes have produced a sales volume ranging from 600 to 800 units per machine per month. The following information is made available to Teguh in evaluating the possible profitability of the operation.

  1. An investment of $45,000 will be required, $9,000 for merchandise and $36,000 for the 20 machines.
  2. The machines have a service life of five years and no salvage value at the end of that period. Depreciation will be computed on the straight-line basis.
  3. The merchandise (candy and soft drinks) retails for an average of 75 cents per unit and will cost Teguh an average of 25 cents per unit.
  4. Owners of the buildings in which the machines are located are paid a commission of 5 cents per unit of candy and soft drinks sold.
  5. One person will be hired to service the machines. The salary will be $1,500 per month.
  6. Other expenses are estimated at $600 per month. These expenses do not vary with the number of units sold.

Required:

a. Determine the unit contribution margin and the break-even volume in units and in dollars per month. (Do not round intermediate calculations. Round "Unit contribution margin" to 2 decimal places.)

c. What sales volume in units and in dollars per month will be necessary to produce an operating income equal to a 30 percent annual return on Teguh's $45,000 investment? (Do not round intermediate calculations.)

d. Teguh is considering offering the building owners a flat rental of $30 per machine per month in lieu of the commission of 5 cents per unit sold. What effect would this change in commission arrangement have on his monthly break-even volume in terms of units? (Do not round intermediate calculations.)

Answer is complete but not entirely correct. $ 0.45 6,000 $4,500 36,333 a. Unit contribution margin Break-even volume in units Break-even volume in dollars c. Sales volume in units Sales volume in dollars 27,250 New break-even volume in units d. 6,600
Answer #1

Solution a:

Contribution margin per unit = Selling price - variable cost = $0.75 - $0.25 - $0.05 = $0.45 per unit

Fixed cost per month = Salary + Other expenses + Depreciation

= $1,500 + $600 + ($36,000/5/12) = $2,700

Breakeven units per month = Fixed cost / Contribution margin per unit = $2,700 / $0.45 = 6000 units

Breakeven volume in dollar = Breakeven units * selling price per unit = 6000*$0.75 = $4,500

Solution c:

Required annual return = ($45,000*30%) = $13,500

Target profit per mont = $13,500/12 = $1,125

Sales volume in units to earn target return = (Fixed cost + Target profit) / Contribution margin per unit

= ($2,700 + $1,125) / $0.45 = 8500 units

Sales volume in dollar = 8500*$0.75 = $6,375

Solution d:

New contribution margin per unit = $0.75 - $0.25 = $0.50 per unit

New fixed cost per month = $2,700 + (20*$30) = $3,300

New breakeven point in units per month = $3,300 / $0.50 = 6600 units